OpenStudy (anonymous):

(absolute value of x minus 5) is less than or equal to 2 How many integers satisfy the inequality above? (A) None (B) One (C) Two (D) Three (E) Five

OpenStudy (anonymous):

sorry I x-5 I less than or equal to 2

hartnn (hartnn):

\(|x-5| \le 2 \) for |a| <=b, we have a<=b or a>=-b so what about \(|x-5| \le 2 \) ??

OpenStudy (anonymous):

how many integers satisfy the equation?

hartnn (hartnn):

lets find out! \(|x-5|\le 2 \implies (x-5)\le2 \quad or \quad (x-5)\ge -2 \) can you soilve them individually ?

hartnn (hartnn):

**solve

OpenStudy (anonymous):

i think so... not so sure...

hartnn (hartnn):

x-5 <= 2 add 5 to both sides, what u get ?

OpenStudy (anonymous):

you get x = 7

hartnn (hartnn):

not x=7 \(x-5 \le 2\) adding 5 to both sides, \(x-5+5 \le 2+5 \implies x \le7 \) you don't loose <= sign

OpenStudy (anonymous):

oh ya i forgot. *facepalm*

hartnn (hartnn):

similarly what u get after adding 5 to both sides of x-5 >= -2 ??

OpenStudy (anonymous):

you would get +3

hartnn (hartnn):

x>=3, right ? so, x>=3 and x<=7 gives you, x=3,4,5,6,7 ----> 5 values

OpenStudy (anonymous):

A great way of solving this problem would be to visually interpret it. When we say: \[\bf |x-5| \le 2\]We're really finding values of 'x' which are within a distance of 2 unit from 5. The only values within 2 units from 5 are 7 and 3. So really, we are looking at all integers within 2 units to the left of 5 on the number line and within 2 units on the right of the number line.Which integers are in that range once you look at the number line? @betterwiththelightsoff99

OpenStudy (anonymous):

oh i understand now.... thank you hartnn

OpenStudy (anonymous):

@genius12 wouldnt that be a little more time-consuming?

OpenStudy (anonymous):

Nope. Once you understand how that works, you can work it out in your head without having to write or type a word. Think about it. We are looking at integers within 2 units of 5. How many are there? 3, 4, 6, 7. 2 units to the left and 2 units to the right. Easy. You can always doing it algebraically like @hartnn but understanding what the |x - 5| really means can be really helpful as you can solve the inequality in your head without much work. @betterwiththelightsoff99

OpenStudy (anonymous):

oh! ok thank you @genius12

OpenStudy (anonymous):

But remember, because its less than or equal to 2, we are finding values that are 2 units or less away from 5. Because its 2 units or less, we include the 3 and 7. If it was just less than 2 units, then the only integers satisfying the inequality would 4 and 6.

OpenStudy (anonymous):

alright got it thanks for the help @genius12

OpenStudy (anonymous):

np